# Translating Flow to Policy via Hindsight Online Imitation

### v1

@@ Acknowledgement
This work is supported by the National Key R&D Program of China (2022ZD0161700), National Natural Science Foundation of China (62176135, 12201341), Shanghai Qi Zhi Institute Innovation Program SQZ202306 and the Tsinghua University Dushi Program.
@@ A.1 Implementation Details of our method
In our pipeline, a crucial step in obtaining flow from video is sampling the task-relevant points within the image coordinates. In simulation experiments, we first acquire segmentation masks directly from the simulation environment and subsequently identify the regions corresponding to the robot end effector and task-relevant objects. Random sampling is then performed within these regions. In real-world experiments, we leverage Grounded SAM 2 (ravi2024sam; ren2024grounded) for segmentation. Prompted with “robot” and “white mouse”, it can segment out the task-relevant part in the image observations. To improve the robustness of the resulting flow data, we apply this sampling strategy multiple times per image during both high-level and low-level training phases.
@@
Owing to the inherent challenges posed by temporally correlated confounders, a single-step policy is likely to encounter significant difficulties in achieving stable training. To address this issue and ensure stable policy training, we have adopted the techniques of action chunking and temporal ensemble, both of which have been widely utilized in prior research (ALOHA; DiffusionPolicy; zhao2024alohaunleashed). Specifically, our flow-guided policy can be formally represented as $\pi_{\theta}(a_{t:t+k}|o_{t},\mathbf{F}_{\text{flow}}(o_{t},\mathbf{p}_{t}))$, as opposed to the conventional single-step policy formulation $\pi_{\theta}(a_{t}|o_{t},\mathbf{F}_{\text{flow}}(o_{t},\mathbf{p}_{t}))$. This implies that at every $k$ steps, the agent generates the subsequent $k$ actions based on the current observation and flow prediction. Furthermore, our temporal ensemble mechanism employs a weighted average over these predictions, utilizing an exponential weighting scheme defined as $w_{i}=\exp(-m\times i)$. In our implementation, the chunk size is consistently set to 5.
@@ Appendix B Experiment Tasks
In this paper, we experiment with 4 tasks based on LIBERO (liu2023libero) and 3 tasks based on ManiSkill3 (taomaniskill3). For each task, we set up a third-person camera and a wrist camera. The image resolution is $128\times 128$. HinFlow samples points both on task-relevant objects and the robot gripper in the third-person camera and sample grid points in the wrist camera. Figure 8 and Figure 9 illustrate the third-party camera view for each task, and Figure 10 shows an example of the wrist camera in LIBERO and ManiSkill environments.
@@
  • Place Book: The robot is required to grasp a book and lay it flat in the left compartment of a desk caddy. We sample 16 points on the book and 16 points on the robot gripper. The horizon is 250.
@@ C.1 Cross-Embodiment Transfer Details
We perform cross-embodiment transfer experiments from Franka to Kinova in LIBERO and from Franka to xArm in ManiSkill. We collect 5 demonstrations for each task via SpaceMouse teleoperation. The detail success rate curve is shown in Figure 11. We can observe that the online imitation may harm the overall performance if the high-level planner doesn’t provide good flow guidance.
@@ C.4.1 UniPi
UniPi (du2023learning) is a hierarchical method capable of effectively learning from action-free video data. It trains a video prediction model to serve as a high-level planner and learns an inverse dynamics model to infer actions from the videos. Since UniPi does not provide an open-source codebase, we follow the implementations in prior work (kolearning; wen2023ATManypointtrajectory) to reproduce this baseline.
@@ C.4.2 PPO
We include Proximal Policy Optimization (PPO) (schulman2017proximal) as a standard reinforcement learning baseline on the ManiSkill tasks. The experiments use the official RGB-based PPO implementation in ManiSkill with the provided dense reward function. As shown in  Figure 15, PPO fails to solve these tasks in 80K environment steps: the success rate curve is below 10% on all three tasks. In contrast, HinFlow significantly improves the success rate within the same online interaction budget.
@@ C.4.3 HuDOR
In this section, we compare our method against HuDOR (guzey2025hudorbridging), which uses object-oriented trajectory rewards to perform online residual reinforcement learning. HuDOR was originally developed for dexterous hand manipulation, so we made several modifications to adapt it to our setting. (1) While HuDOR uses human-hand-retargeted trajectories as the base policy, we replace that with the same pretrained policy used by our method. (2) To learn a residual policy via reinforcement learning, HuDOR relies on object-centric trajectories extracted from human videos for reward shaping. As an alternative, at each simulation reset, we generate an action-free success video from the current initial state using a scripted policy and provide it to HuDOR for reward computation. This is a form of privileged information that HinFlow does not receive. (3) We follow HuDOR’s implementation for the residual policy architecture and RL algorithm. The residual policy’s input is a vector concatenating the base policy action, object-centric translational and rotational features, the robot’s proprioceptive state, and the base policy’s latent visual feature (which encodes observations from two cameras). The residual scale is 1.0 for the gripper action dimension and 0.05 for all other action dimensions. (4) Since the wrist-mounted camera cannot consistently observe the object, in both reward shaping and residual policy inputs, we only use the object-centric translational and rotational features from the third-person camera. (5) In addition, on top of the RL loss, we add an auxiliary behavior cloning loss computed from the action-labeled demonstrations, which we find improves the performance of the residual policy.

### v2

@@ Acknowledgement
This work is supported by Shanghai Qi Zhi Institute & Spirit AI Innovation Program and the Tsinghua University Dushi Program.
@@ A.1 Implementation Details of our method
In our pipeline, a crucial step in obtaining flow from video is sampling the task-relevant points within the image coordinates. In simulation experiments, we first acquire segmentation masks directly from the simulation environment and subsequently identify the regions corresponding to the robot end effector and task-relevant objects. Random sampling is then performed within these regions. In real-world experiments, we leverage Grounded SAM 2 (Ravi et al., 2024; Ren et al., 2024) for segmentation. Prompted with “robot” and “white mouse”, it can segment out the task-relevant part in the image observations. To improve the robustness of the resulting flow data, we apply this sampling strategy multiple times per image during both high-level and low-level training phases.
@@
Owing to the inherent challenges posed by temporally correlated confounders, a single-step policy is likely to encounter significant difficulties in achieving stable training. To address this issue and ensure stable policy training, we have adopted the techniques of action chunking and temporal ensemble, both of which have been widely utilized in prior research (Zhao et al., 2023; Chi et al., 2023; Zhao et al., 2024). Specifically, our flow-guided policy can be formally represented as $\pi_{\theta}(a_{t:t+k}|o_{t},\mathbf{F}_{\text{flow}}(o_{t},\mathbf{p}_{t}))$, as opposed to the conventional single-step policy formulation $\pi_{\theta}(a_{t}|o_{t},\mathbf{F}_{\text{flow}}(o_{t},\mathbf{p}_{t}))$. This implies that at every $k$ steps, the agent generates the subsequent $k$ actions based on the current observation and flow prediction. Furthermore, our temporal ensemble mechanism employs a weighted average over these predictions, utilizing an exponential weighting scheme defined as $w_{i}=\exp(-m\times i)$. In our implementation, the chunk size is consistently set to 5.
@@ Appendix B Experiment Tasks
In this paper, we experiment with 4 tasks based on LIBERO (Liu et al., 2023) and 3 tasks based on ManiSkill3 (Tao et al., 2025). For each task, we set up a third-person camera and a wrist camera. The image resolution is $128\times 128$. HinFlow samples points both on task-relevant objects and the robot gripper in the third-person camera and sample grid points in the wrist camera. Figure 8 and Figure 9 illustrate the third-party camera view for each task, and Figure 10 shows an example of the wrist camera in LIBERO and ManiSkill environments.
@@
  • Place Book: The robot is required to grasp a book and lay it flat in the left compartment of a desk caddy. We sample 16 points on the book and 16 points on the robot gripper. The horizon is 200.
@@ C.1 Cross-Embodiment Transfer Details
We perform cross-embodiment transfer experiments from Franka to Kinova in LIBERO and from Franka to xArm in ManiSkill. We collect 5 demonstrations for each task. The detail success rate curve is shown in Figure 11. We can observe that the online imitation may harm the overall performance if the high-level planner doesn’t provide good flow guidance.
@@ C.4.1 UniPi
UniPi (Du et al., 2023) is a hierarchical method capable of effectively learning from action-free video data. It trains a video prediction model to serve as a high-level planner and learns an inverse dynamics model to infer actions from the videos. Since UniPi does not provide an open-source codebase, we follow the implementations in prior work (Ko et al., 2023; Wen et al., 2023) to reproduce this baseline.
@@ C.4.2 PPO
We include Proximal Policy Optimization (PPO) (Schulman et al., 2017) as a standard reinforcement learning baseline on the ManiSkill tasks. The experiments use the official RGB-based PPO implementation in ManiSkill with the provided dense reward function. As shown in  Figure 15, PPO fails to solve these tasks in 80K environment steps: the success rate curve is below 10% on all three tasks. In contrast, HinFlow significantly improves the success rate within the same online interaction budget.
@@ C.4.3 HuDOR
In this section, we compare our method against HuDOR (Guzey et al., 2025), which uses object-oriented trajectory rewards to perform online residual reinforcement learning. HuDOR was originally developed for dexterous hand manipulation, so we made several modifications to adapt it to our setting. (1) While HuDOR uses human-hand-retargeted trajectories as the base policy, we replace that with the same pretrained policy used by our method. (2) To learn a residual policy via reinforcement learning, HuDOR relies on object-centric trajectories extracted from human videos for reward shaping. As an alternative, at each simulation reset, we generate an action-free success video from the current initial state using a scripted policy and provide it to HuDOR for reward computation. This is a form of privileged information that HinFlow does not receive. (3) We follow HuDOR’s implementation for the residual policy architecture and RL algorithm. The residual policy’s input is a vector concatenating the base policy action, object-centric translational and rotational features, the robot’s proprioceptive state, and the base policy’s latent visual feature (which encodes observations from two cameras). The residual scale is 1.0 for the gripper action dimension and 0.05 for all other action dimensions. (4) Since the wrist-mounted camera cannot consistently observe the object, in both reward shaping and residual policy inputs, we only use the object-centric translational and rotational features from the third-person camera. (5) In addition, on top of the RL loss, we add an auxiliary behavior cloning loss computed from the action-labeled demonstrations, which we find improves the performance of the residual policy.